The owner of Genuine Subs, Inc., hopes to expand the present operation by adding one new outlet. She has studied three locations. Each would have the same labor and materials costs (food, serving containers, napkins, etc.) of $2.20 per sandwich. Sandwiches sell for $3.00 each in all locations. Rent and equipment costs would be $5,550 per month for location A, $5,775 per month for location B, and $6,025 per month for location C.

a. Determine the volume necessary at each location to realize a monthly profit of $10,750. (Do not round intermediate calculations. Round your answer to the nearest whole number.) Location Monthly Volume

A $ _____

B $ _____

C $ _____

b-1. If expected sales at A, B, and C are 24,750 per month, 21,750 per month, and 23,750 per month, respectively, calculate the profit of the each locations? (Omit the "$" sign in your response.)

Location Monthly Profits

A $_____

B $_____

C $ _____

b-2. Which location would yield the greatest profits?

Location C

Location B

Location A
Answer #1

a. Profit = Revenue - Expense

For location A, let the volume be X, then

10750 = (3*X) - (5550 + 2.2*X)

0.8*X = 16300

X = 20375

For location B, let the volume be Y, then

10750 = 3*Y - (2.2*Y + 5775)

Y = 20656

For location C, let the volume be Z, then

10750 = 3*Z - (2.2*Z + 6025)

Z = 20969

b-1. Profit for A

= 3*24750 - (2.2*24750 + 5550)

= $14250

Profit for B

= 3*21750 - (2.2*21750 + 5775)

= $11625

Profit for C

= 3*23750 - (2.2*23750+6025)

= $12975

b-2. location A

Based on the profit amount.